OpenStudy (anonymous):

If one of the roots of the quadratic equation 3x^2+kx+6=0 is 1.5, then what is the value of k?

OpenStudy (anonymous):

that was wrong

OpenStudy (anonymous):

but in fact we can find it a couple different ways. you know that the product of the roots is \[\frac{c}{a}\] and the sum is \[-\frac{b}{a}\]

OpenStudy (anonymous):

(3x+a)(x-1.5) 3x^2+ax-4.5x-1.5a 1.5a=6 a=-4 k=-8.5

OpenStudy (anonymous):

Thank you so much!